# Index Terms: Time-Series Analysis, Embedded AI, Tiny Transformers, Model Quantization, FPGAs AccelerationI Introduction

### v4

@@ Abstract
Transformer-based models have shown strong performance across diverse time-series tasks, but their deployment on resource-constrained devices remains challenging due to high memory and computational demand. While prior work targeting Microcontroller Units (MCUs) has explored hardware-specific optimizations, such approaches are often task-specific and limited to 8-bit fixed-point precision. Field-Programmable Gate Arrays (FPGAs) offer greater flexibility, enabling fine-grained control over data precision and architecture. However, existing FPGA-based deployments of Transformers for time-series analysis typically focus on high-density platforms with manual configuration. This paper presents a unified and fully automated deployment framework for Tiny Transformers on embedded FPGAs. Our framework supports a compact encoder-only Transformer architecture across three representative time-series tasks (forecasting, classification, and anomaly detection). It combines quantization-aware training (down to 4 bits), hardware-aware hyperparameter search using Optuna, and automatic VHDL generation for seamless deployment. We evaluate our framework on six public datasets across two embedded FPGA platforms. Results show that our framework produces integer-only, task-specific Transformer accelerators achieving as low as 0.033 mJ per inference with millisecond latency on AMD Spartan-7, while also providing insights into deployment feasibility on Lattice iCE40. All source code will be released in the GitHub repository111https://github.com/tianheng-ling/TinyTransformer4TS.
Index Terms:
Time-Series Analysis, Embedded AI, Tiny Transformers, Model Quantization, FPGAs AccelerationI Introduction
@@ III-A Overview of Deployment Framework
As shown in Figure 1, our framework begins by designing and configuring a QuantTransformer model (❷) using quantization-compatible layer or module from an open-source ElasticAI.Creator222https://github.com/es-ude/elastic-ai.creator/tree/add-linear-quantization library [18]. QAT is applied to calculate quantization parameters for later enabling integer-only inference. Once training converges, the model is exported to an integer format (❸) suitable for hardware mapping. In the VHDL generation phase (❹), each quantized layer is translated to a RTL module with parameterized bitwidths and tensor dimensions. The design is first functionally verified using GHDL RTL simulation (❺) and then synthesized with Vivado (for AMD Spartan-7) or Radiant (for Lattice iCE40) (❻), generating post-synthesis reports on logic utilization, timing, and estimated power consumption. Inspired by hardware-aware optimization in LOTTA [19], we integrate an Optuna-driven hyperparameter search loop (❼). Each candidate configuration undergoes QAT, simulation, synthesis, and deployability filtering. Only configurations that pass hardware constraints proceed to final validation. To establish a performance baseline, each selected configuration is also evaluated using a PyTorch-based FloatTransformer model (❽) trained and validated on the same dataset in full-precision (FP32). The final quantized models are then deployed on real hardware for end-to-end testing and performance validation (❾). We target two embedded FPGA platforms: ElasticNode V5 (AMD Spartan-7 XC7S15 @100 MHz, 8,000 LUTs, 20 DSPs, 10 BRAMs), and ElasticNode V5 SE (Lattice iCE40UP5K @16 MHz, 5,280 LUTs, 8 DSPs, 30 ERBs). Both platforms integrate an RP2040 Cortex-M0+ MCU coordinating data acquisition and managing inference requests to the FPGA [20].
@@ IV-A Experimental Setup
Task | Datasets | Configuration | Accuracy Metrics‡ | LUTs (%) | BRAMs (%) | DSPs (%) | Energy (mJ) | Power (mW)∗ | Latency (ms)∗∗ | |||||
b | bs | lr ( | params† | FP32 | INT | |||||||||
Forecasting | PeMS | 6 | 208 | 6.357 | 16 | 3329 | 0.160 | 0.194 | 46.60 | 100 | 90 | 0.078 | 65.0 | 1.203 |
AirU | 8 | 32 | 5.025 | 8 | 897 | 4.129 | 4.853 | 53.48 | 95 | 100 | 0.036 | 64.0 | 0.570 | |
Classification | HCIHAR | 8 | 240 | 4.618 | 8 | 1006 | 0.858 | 0.824 | 53.64 | 100 | 90 | 0.067 | 65.0 | 1.034 |
WISDM | 6 | 48 | 1.257 | 40 | 20126 | 0.838 | 0.839 | 96.94 | 100 | 100 | 0.855 | 71.0 | 12.04 | |
Anomaly Detection | ALFA | 4 | 192 | 1.299 | 8 | 1106 | 0.923 | 0.889 | 35.55 | 100 | 65 | 0.033 | 62.0 | 0.527 |
SKAB | 6 | 96 | 5.064 | 24 | 7465 | 0.766 | 0.765 | 58.95 | 100 | 95 | 0.154 | 68.0 | 2.261 | |
†The number of model parameters. | ||||||||||||||
‡Test RMSE for forecasting task, Test Accuracy for classification task, and Test F1 score for anomaly detection task. | ||||||||||||||
∗ Power was estimated at a temperature of 28.0 ∘C with a power deviation of 5.8% compared to actual hardware measurements. | ||||||||||||||
∗∗ The latency obtained from GHDL compared to actual hardware measurements varies by about 2%. |
@@ IV-B Time-Series Forecasting
We begin by evaluating our framework on two datasets using a 24-step sliding window for single-step ahead forecasting. PeMS333https://doi.org/10.5281/zenodo.3939793 is a univariate traffic flow dataset with 5-minute readings from over 11,000 sensors. Following [16], we extract data from sensor 4192. AirU444https://dx.doi.org/10.21227/aeh2-a413 is a multivariate air quality dataset containing 19,380 samples across seven features. Based on [7], we retain 15,258 valid samples and use ozone concentration as the prediction target, applying Min-Max normalization across all features. During Optuna search, models are optimized using the Mean Squared Error (MSE) loss, and RMSE (calculated on denormalized targets and predictions) is used for final evaluation to ensure fair comparison with [16].
@@ IV-C Time-Series Classification
We next evaluate our framework on two human activity recognition datasets to assess its suitability for classification tasks. UCIHAR555https://github.com/arijitiiest/UCI-Human-Activity-Recognition dataset comprises 9 sensor signals (accelerometer and gyroscope, 3-axis each) sampled at 50 Hz. To reduce computational overhead, we apply a 4 downsampling, resulting in 32-step windows for 6-class classification. WISDM666https://github.com/bartkowiaktomasz/har-wisdm-lstm-rnns dataset provides motion data with three channels. We downsample the input by 4 to obtain 50-step sequences, and limit experiments to 1600 samples per class due to constrained training resources. During Optuna search, models are optimized using Cross-entropy loss and selected based on macro accuracy. Figure 4 presents the Pareto fronts for both datasets, where red markers denote configurations that balance validation accuracy and energy on the XC7S15 FPGA.
@@ IV-D Time-Series Anomaly Detection
We conclude our evaluation with two threshold-based anomaly detection datasets. ALFA777https://github.com/castacks/alfa-dataset dataset captures multivariate UAV telemetry with 17 channels under fault injection (e.g., motor failure, sensor drift). Following [22], we predict 10 status variables one step ahead and apply smoothing-based residual thresholding, sweeping to select thresholds minimizing deviation from peak residuals. SKAB888https://github.com/waico/SKAB dataset contains 8-channel sensor readings from a water pump system under synthetic anomalies. We forecast the next flow rate and determine detection thresholds () via a percentile-based rule on absolute residuals(). All models are trained using MSE loss, and inference is performed by comparing predicted residuals against fixed thresholds computed on the validation set.

### v5

# I Introduction

@@ Abstract
Transformer-based models have shown strong performance across diverse time-series tasks, but their deployment on resource-constrained devices remains challenging due to high memory and computational demand. While prior work targeting Microcontroller Units (MCUs) has explored hardware-specific optimizations, such approaches are often task-specific and limited to 8-bit fixed-point precision. Field-Programmable Gate Arrays (FPGAs) offer greater flexibility, enabling fine-grained control over data precision and architecture. However, existing FPGA-based deployments of Transformers for time-series analysis typically focus on high-density platforms with manual configuration. This paper presents a unified and fully automated deployment framework for Tiny Transformers on embedded FPGAs. Our framework supports a compact encoder-only Transformer architecture across three representative time-series tasks (forecasting, classification, and anomaly detection). It combines quantization-aware training (down to 4 bits), hardware-aware hyperparameter search using Optuna, and automatic VHDL generation for seamless deployment. We evaluate our framework on six public datasets across two embedded FPGA platforms. Results show that our framework produces integer-only, task-specific Transformer accelerators achieving as low as 0.033 mJ per inference with millisecond latency on AMD Spartan-7, while also providing insights into deployment feasibility on Lattice iCE40. All source code will be released at: https://github.com/tianheng-ling/TinyTransformer4TS.
I Introduction
@@ III-A Overview of Deployment Framework
As shown in Figure 1, our framework begins by designing and configuring a QuantTransformer model (❷) using quantization-compatible layer or module from an open-source ElasticAI.Creator111https://github.com/es-ude/elastic-ai.creator/tree/add-linear-quantization library [18]. QAT is applied to calculate quantization parameters for later enabling integer-only inference. Once training converges, the model is exported to an integer format (❸) suitable for hardware mapping. In the VHDL generation phase (❹), each quantized layer is translated to a RTL module with parameterized bitwidths and tensor dimensions. The design is first functionally verified using GHDL RTL simulation (❺) and then synthesized with Vivado (for AMD Spartan-7) or Radiant (for Lattice iCE40) (❻), generating post-synthesis reports on logic utilization, timing, and estimated power consumption. Inspired by hardware-aware optimization in LOTTA [19], we integrate an Optuna-driven hyperparameter search loop (❼). Each candidate configuration undergoes QAT, simulation, synthesis, and deployability filtering. Only configurations that pass hardware constraints proceed to final validation. To establish a performance baseline, each selected configuration is also evaluated using a PyTorch-based FloatTransformer model (❽) trained and validated on the same dataset in full-precision (FP32). The final quantized models are then deployed on real hardware for end-to-end testing and performance validation (❾). We target two embedded FPGA platforms: ElasticNode V5 (AMD Spartan-7 XC7S15 @100 MHz, 8,000 LUTs, 20 DSPs, 10 BRAMs), and ElasticNode V5 SE (Lattice iCE40UP5K @16 MHz, 5,280 LUTs, 8 DSPs, 30 ERBs). Both platforms integrate an RP2040 Cortex-M0+ MCU coordinating data acquisition and managing inference requests to the FPGA [20].
@@ IV-A Experimental Setup
| Task | Datasets | Configuration | Accuracy Metrics‡ |
| b | bs | lr ( | params† | FP32 | INT | |||||||||||||||||||||
| Forecasting | PeMS | 6 | 208 | 6.357 | 16 | 3329 | 0.160 | 0.194 | 46.60 | 100 | 90 | 0.078 | 65.0 | 1.203 | ||||||||||||
| AirU | 8 | 32 | 5.025 | 8 | 897 | 4.129 | 4.853 | 53.48 | 95 | 100 | 0.036 | 64.0 | 0.570 | |||||||||||||
| Classification | UCIHAR | 8 | 240 | 4.618 | 8 | 1006 | 0.858 | 0.824 | 53.64 | 100 | 90 | 0.067 | 65.0 | 1.034 | ||||||||||||
| WISDM | 6 | 48 | 1.257 | 40 | 20126 | 0.838 | 0.839 | 96.94 | 100 | 100 | 0.855 | 71.0 | 12.04 | |||||||||||||
ALFA | 4 | 192 | 1.299 | 8 | 1106 | 0.923 | 0.889 | 35.55 | 100 | 65 | 0.033 | 62.0 | 0.527 | ||||||||||||
| SKAB | 6 | 96 | 5.064 | 24 | 7465 | 0.766 | 0.765 | 58.95 | 100 | 95 | 0.154 | 68.0 | 2.261 | |||||||||||||
| †The number of model parameters. | ||||||||||||||||||||||||||
| ‡Test RMSE for forecasting task, Test Accuracy for classification task, and Test F1 score for anomaly detection task. | ||||||||||||||||||||||||||
| ∗ Power was estimated at a temperature of 28.0 ∘C with a power deviation of 5.8% compared to actual hardware measurements. | ||||||||||||||||||||||||||
| ∗∗ The latency obtained from GHDL compared to actual hardware measurements varies by about 2%. | ||||||||||||||||||||||||||
@@ IV-B Time-Series Forecasting
We begin by evaluating our framework on two datasets using a 24-step sliding window for single-step ahead forecasting. PeMS222https://doi.org/10.5281/zenodo.3939793 is a univariate traffic flow dataset with 5-minute readings from over 11,000 sensors. Following [16], we extract data from sensor 4192. AirU333https://dx.doi.org/10.21227/aeh2-a413 is a multivariate air quality dataset containing 19,380 samples across seven features. Based on [7], we retain 15,258 valid samples and use ozone concentration as the prediction target, applying Min-Max normalization across all features. During Optuna search, models are optimized using the Mean Squared Error (MSE) loss, and RMSE (calculated on denormalized targets and predictions) is used for final evaluation to ensure fair comparison with [16].
@@ IV-C Time-Series Classification
We next evaluate our framework on two human activity recognition datasets to assess its suitability for classification tasks. UCIHAR444https://github.com/arijitiiest/UCI-Human-Activity-Recognition dataset comprises 9 sensor signals (accelerometer and gyroscope, 3-axis each) sampled at 50 Hz. To reduce computational overhead, we apply a 4 downsampling, resulting in 32-step windows for 6-class classification. WISDM555https://github.com/bartkowiaktomasz/har-wisdm-lstm-rnns dataset provides motion data with three channels. We downsample the input by 4 to obtain 50-step sequences, and limit experiments to 1600 samples per class due to constrained training resources. During Optuna search, models are optimized using Cross-entropy loss and selected based on macro accuracy. Figure 4 presents the Pareto fronts for both datasets, where red markers denote configurations that balance validation accuracy and energy on the XC7S15 FPGA.
@@ IV-D Time-Series Anomaly Detection
We conclude our evaluation with two threshold-based anomaly detection datasets. ALFA666https://github.com/castacks/alfa-dataset dataset captures multivariate UAV telemetry with 17 channels under fault injection (e.g., motor failure, sensor drift). Following [22], we predict 10 status variables one step ahead and apply smoothing-based residual thresholding, sweeping to select thresholds minimizing deviation from peak residuals. SKAB777https://github.com/waico/SKAB dataset contains 8-channel sensor readings from a water pump system under synthetic anomalies. We forecast the next flow rate and determine detection thresholds () via a percentile-based rule on absolute residuals(). All models are trained using MSE loss, and inference is performed by comparing predicted residuals against fixed thresholds computed on the validation set.